P4(s)+10Cl2(g)→4PCl5(g)

Now consider a situation in which 29.0 g of P4 is added to 55.0 g of Cl2, and a chemical reaction occurs. To identify the limiting reactant, you will need to perform two separate calculations:

1. Calculate the number of moles of PCl5 that can be produced from 29.0 g of P4 (and excess Cl2).

2. Calculate the number of moles of PCl5 that can be produced from 55.0 g of Cl2 (and excess P4).

3. Then, compare the two values. The reactant that produces the smaller amount of product is the limiting reactant.
Answer #1

1)

Molar mass of P4 = 123.88 g/mol

mass of P4 = 29 g

mol of P4 = (mass)/(molar mass)

= 29/1.239*10^2

= 0.2341 mol

Balanced chemical equation is:

P4(s)+10Cl2(g)→4PCl5(g)

According to balanced equation

mol of PCl5 formed = (4/1)* moles of P4

= (4/1)*0.2341

= 0.9364 mol

Answer: 0.936 mol

2)

Molar mass of Cl2 = 70.9 g/mol

mass of Cl2 = 55 g

mol of Cl2 = (mass)/(molar mass)

= 55/70.9

= 0.7757 mol

Balanced chemical equation is:

P4(s)+10Cl2(g)→4PCl5(g)

According to balanced equation

mol of PCl5 formed = (4/10)* moles of Cl2

= (4/10)*0.7757

= 0.3103 mol

Answer: 0.310 mol

3)

Cl2 produces smaller mol of PCl5

So, Cl2 is limiting reagent

Answer: Cl2